# Two-loop anomalous dimensions for baryon-number-violating operators in SMEFT

### v1

@@ 1 Introduction
In these cases, BNV usually occurs are high energies to suppress its effects at low energies and ensure a sufficiently long proton lifetime to satisfy the experimental constraints. In such a setup, the BNV effects can be parametrized within effective field theories (EFTs) via higher-dimension operators, starting at dimension 6. As the BSM scale is usually much higher than the electroweak scale, one can use the SM Effective Field Theory (SMEFT) with explicit $SU(2)_{L}$ invariance to reduce the number of operators Buchmuller:1985jz ; Grzadkowski:2010es , see Refs. Crivellin:2023ter ; Beneito:2023xbk ; Gargalionis:2024nij ; Beneke:2024hox ; Gisbert:2024sjw ; Broussard:2025opd for recent work. For the evolution of the Wilson coefficients below the electroweak scale, the SMEFT should be matched to the low-energy EFT (LEFT). This EFT approach not only simplifies the calculations but also improves the precision since the potentially huge scale separation between the BNV scale and the hadronic scale mandates the resummation of large logarithms.
@@
The one-loop renormalization-group equations (RGEs) of these two EFTs at dimension 6 Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga ; Alonso:2014zka ; Jenkins:2017dyc , together with the tree-level matching of the SMEFT on LEFT Jenkins:2017jig can be used for computations at leading-logarithmic (LL) accuracy. For next-to-leading-log (NLL) accuracy the one-loop matching between SMEFT and LEFT Dekens:2019ept , the two-loop RGEs Buras:1989xd ; Buras:1991jm ; Buras:1992tc ; Ciuchini:1993vr ; Ciuchini:1993fk ; Buchalla:1995vs ; Chetyrkin:1997gb ; Buras:2000if ; Gorbahn:2016uoy ; Panico:2018hal ; deVries:2019nsu ; Bern:2020ikv ; Aebischer:2022anv ; Fuentes-Martin:2022vvu ; Aebischer:2023djt ; Jenkins:2023rtg ; Jenkins:2023bls ; Naterop:2023dek ; DiNoi:2023ygk ; Fuentes-Martin:2023ljp ; Morell:2024aml ; Aebischer:2024xnf ; Manohar:2024xbh ; DiNoi:2024ajj ; Born:2024mgz ; Naterop:2024cfx ; Fuentes-Martin:2024agf ; Aebischer:2025hsx ; Duhr:2025zqw ; Haisch:2025lvd ; Zhang:2025ywe ; Naterop:2025lzc ; Naterop:2025cwg ; Duhr:2025yor , as well as the calculation of all observables at next-to-leading order (NLO) are necessary. Furthermore, the one-loop matching of UV models on the SMEFT at NLO is needed, which has been automated to a large extent Carmona:2021xtq ; Fuentes-Martin:2022jrf ; Fuentes-Martin:2023ljp ; Aebischer:2023nnv ; Thomsen:2024abg .
@@
For the BNV sector at dimension 6, the one-loop RGEs were calculated in Refs. Abbott:1980zj ; Alonso:2014zka ; Jenkins:2017dyc . In the LEFT, the two-loop RGEs are already available in multiple schemes Aebischer:2025hsx ; Naterop:2025lzc . [^1] The disagreement between these two references is expected to disappear with an upcoming revision of Ref. Aebischer:2025hsx . In the following, we rely on the results of Ref. Naterop:2025lzc . In the present article, we complete the two-loop RGE calculation for the BNV SMEFT operators at dimension 6. In addition to $SU(3)_{c}$ and $U(1)_{Y}$ gauge contributions as in the LEFT, our calculation includes the $SU(2)_{L}$ gauge and Yukawa interactions, as well as mixed gauge–Yukawa two-loop contributions and the Higgs self-interactions. In addition, we calculate the one-loop correction to the matching on SMEFT of the $S_{1}$ leptoquark, which can generate the Wilson coefficients of all four gauge-invariant operators. Therefore, our work completes the NLL EFT framework for the BNV sector at dimension 6, enabling refined analyses of the constraints from proton decay on heavy BNV new physics if combined with the NLO matching corrections Gracey:2012gx to non-perturbative determinations of the hadronic matrix elements Yoo:2021gql .
@@ 3.1 Dimensional regularization
As in the LEFT Naterop:2025lzc ; Aebischer:2025hsx , the two-loop renormalization of the BNV sector of the SMEFT does not involve any ill-defined $\gamma_{5}$-odd traces. Therefore, we perform the calculation in naive dimensional regularization (NDR) with anti-commuting $\gamma_{5}$, avoiding the issue of spurious symmetry breaking of the ’t Hooft and Veltman scheme and the need for finite symmetry-restoring renormalizations tHooft:1972tcz ; Breitenlohner:1977hr .
@@ 6.3 Cancellation of scheme and scale dependences
In the following, we check the cancellation of scheme and scale dependences between the two-loop RGEs and the one-loop matching. To this end, we relate the renormalized $\overline{\text{MS}}$ LQ mass $M$ to the pole mass $M_{\mathrm{OS}}$,
@@
where $\mu$ is the matching scale, $C_{F}=4/3$, $\alpha_{1}=g_{1}^{2}/(4\pi)$, $\alpha_{s}=g_{3}^{2}/(4\pi)$, $\langle\cdot\rangle$ denotes the trace in flavor space, and we neglected the fermion masses, i.e., we compute the pole mass in the unbroken phase. Note that due to the fermionic decay channels, the LQ acquires a width, i.e., the pole mass gets an imaginary part. We included the effects due to higher powers of the LQ Yukawa couplings just for completeness, although they are not relevant for our checks of the cancellation of scale and scheme dependences. Since the pole mass is scale independent, the one-loop RGE for the running $\overline{\text{MS}}$ mass can be read off Eq. (6.6)
$\dot{M}]_{1}=M\begin{aligned} &\Big(-3g_{1}^{2}y_{S_{1}}^{2}-4g_{3}^{2}+\langle Y^{RR}Y^{RR\dagger}\rangle+2\langle Y^{LL}Y^{LL\dagger}\rangle\\ &\quad+2\langle Y^{Q,RR}Y^{Q,RR\dagger}\rangle+8\langle Y^{Q,LL}Y^{Q,LL\dagger}\rangle\Big)\,.\end{aligned}$ (6.7)
For the LQ Yukawa couplings, we use the $\overline{\text{MS}}$ scheme. Their scale dependence at one loop is given by
@@
$\quad+Y^{Q,RR}\left(\langle Y^{RR}Y^{RR\dagger}\rangle+2\langle Y^{LL}Y^{LL\dagger}\rangle+2\langle Y^{Q,RR}Y^{Q,RR\dagger}\rangle+8\langle Y^{Q,LL}Y^{Q,LL\dagger}\rangle\right)\,$, (6.8)
where the corrections due to higher powers of the LQ Yukawa couplings are again included for completeness but are not needed for the check of scheme- and scale-dependence cancellation. For a phenomenological application, the renormalized $\overline{\text{MS}}$ LQ Yukawa couplings should be related to the physical LQ decay widths at NLO, i.e., including virtual and real corrections. Here, we do not compute the LQ decays at NLO but simply assume that the $\overline{\text{MS}}$ couplings have been determined at the fixed scale $\mu=M_{\mathrm{OS}}$.
We now illustrate the cancellation of scheme and scale dependences, restricting the discussion for simplicity to the gauge interactions. With the full two-loop RGE results provided as supplementary material, the analysis can be extended to the Yukawa interactions. We start by defining observables at the scale of the LQ mass $\mu=M_{\mathrm{OS}}\approx M$, given by the LQ pole mass $M_{\mathrm{OS}}$ itself and the $\overline{\text{MS}}$ LQ Yukawa couplings $Y^{LL}(M_{\mathrm{OS}})$, $Y^{RR}(M_{\mathrm{OS}})$, $Y^{Q,LL}(M_{\mathrm{OS}})$, and $Y^{Q,RR}(M_{\mathrm{OS}})$. We then relate the $\overline{\text{MS}}$ parameters at $\mu=M_{\mathrm{OS}}$ to the ones at the matching scale $\mu=\mu_{M}$, to NLO or LL accuracy. Since $\mu_{M}\approx M_{\mathrm{OS}}$, this relation does not involve large logarithms and the difference between a fixed-order NLO relation and a resummation is immaterial. We then determine the SMEFT Wilson coefficients at the scale $\mu=\mu_{M}$ using the one-loop matching. We subsequently apply the two-loop SMEFT RGEs to determine the Wilson coefficients at some low scale $\mu_{L}$, resumming large logarithms to NLL accuracy. When using the flavor decompositions that have diagonal gauge corrections, we can conveniently express the results in terms of the tree-level expressions Eqs. (6.1) and (6.1) with renormalized Yukawa couplings $Y(M_{\mathrm{OS}})$ and pole mass $M_{\mathrm{OS}}$.
For an analytic check of the matching scale cancellation, we expand the solution of the RGEs in powers of the $\overline{\text{MS}}$ gauge couplings at $\mu=M_{\mathrm{OS}}$. At LL accuracy, re-expanded to NLO, this leads to
@@
$\frac{C_{duq\ell}(\mu_{L})}{C_{duq\ell}^{\mathrm{tree}}}\Big|^{\mathrm{NLL}}_{\mathrm{NNLO}}$ $=1-\left(\frac{11\alpha_{1}}{24\pi}+\frac{9\alpha_{2}}{8\pi}+\frac{\alpha_{s}}{\pi}\right)\log\!\left(\frac{\mu_{L}}{M_{\mathrm{OS}}}\right)$
$\quad+\left(-\frac{781\alpha_{1}^{2}}{1152\pi^{2}}+\frac{195\alpha_{2}^{2}}{128\pi^{2}}+\frac{9\alpha_{s}^{2}}{4\pi^{2}}+\frac{33\alpha_{1}\alpha_{2}}{64\pi^{2}}+\frac{11\alpha_{1}\alpha_{s}}{24\pi^{2}}+\frac{9\alpha_{2}\alpha_{s}}{8\pi^{2}}\right)\log^{2}\!\left(\frac{\mu_{L}}{M_{\mathrm{OS}}}\right)$
$\quad+\frac{(2-3a_{\mathrm{ev}})\alpha_{1}}{18\pi}+\frac{7\alpha_{s}}{3\pi}$
$\quad+\bigg(\begin{aligned} &\frac{(1817-3408a_{\mathrm{ev}})\alpha_{1}^{2}}{6912\pi^{2}}-\frac{541\alpha_{2}^{2}}{256\pi^{2}}-\frac{29\alpha_{s}^{2}}{8\pi^{2}}\\ &-\frac{(13-24a_{\mathrm{ev}})\alpha_{1}\alpha_{2}}{128\pi^{2}}-\frac{(145-24a_{\mathrm{ev}})\alpha_{1}\alpha_{s}}{144\pi^{2}}-\frac{33\alpha_{2}\alpha_{s}}{16\pi^{2}}\bigg)\log\!\left(\frac{\mu_{L}}{M_{\mathrm{OS}}}\right)\end{aligned}$
$\quad\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}+\left(-\frac{1379\alpha_{1}^{2}}{2304\pi^{2}}+\frac{541\alpha_{2}^{2}}{256\pi^{2}}+\frac{31\alpha_{s}^{2}}{24\pi^{2}}-\frac{3\alpha_{1}\alpha_{2}}{128\pi^{2}}-\frac{25\alpha_{1}\alpha_{s}}{144\pi^{2}}-\frac{9\alpha_{2}\alpha_{s}}{16\pi^{2}}\right)\log\!\left(\frac{\mu_{M}}{M_{\mathrm{OS}}}\right)$
$\quad+\ldots\,$, (6.10)
where the couplings are evaluated at $\mu=M_{\mathrm{OS}}$ and we dropped non-logarithmic two-loop terms (which come from finite one-loop matching contributions and the one-loop relation to observables and are beyond the order considered here). The NLL result allows us to make a few observations. The first two lines in Eq. (6.3) contain the LL terms, while the next three lines are NLL terms (both truncated at NNLO). While the leading logs are scheme independent, the NLL terms contain scheme dependences, in particular a dependence on $a_{\mathrm{ev}}$. The reason is that the Wilson coefficient itself is a scheme-dependent quantity beyond tree-level. The last line contains a logarithm of the matching scale over the LQ mass, which is not a large logarithm. These terms (in gray) are beyond the accuracy of our calculation, as we did not compute the renormalization of the LQ model at two loops. We find that the dependence on the matching scale $\mu_{M}$ is shifted to NNLO as expected, i.e., all NLO matching-scale dependences cancel. Regarding the scheme dependence, we find that the matching-scale-dependent terms do not depend on $a_{\mathrm{ev}}$: we observe a cancellation of such $a_{\mathrm{ev}}$-dependent terms between the two-loop RGE and the one-loop matching. The remaining $a_{\mathrm{ev}}$-dependence in Eq. (6.3) will cancel instead at the low scale $\mu_{L}$, when combining the two-loop RGE with a NLO matrix-element or matching calculation. The same structure is found for the NLL results of the other Wilson coefficients, which are included in the supplementary material.
@@
$C(\mu=M_{\mathrm{OS}},\mu_{M})/C(\mu=M_{\mathrm{OS}},\mu_{M}=M_{\mathrm{OS}})$
$C_{duq\ell}LLNLL$$C_{qque}$$C_{duue}^{(+)}LLNLL$$C_{duue}^{(-)}$$\mu_{M}/M_{\mathrm{OS}}$$S_{qqq\ell}LLNLL$$M_{qqq\ell}$
Figure 1: Comparison of the LL and NLL dependence of the Wilson coefficients on the matching scale $\mu_{M}$, normalized to a matching scale of $\mu_{M}=M_{\mathrm{OS}}$, which is chosen at the unification scale, $M_{\mathrm{OS}}=6.5\times 10^{15}\,\text{GeV}$. We only include the contribution of the gauge interactions.
The reduction of the matching-scale dependence from NLO to NNLO is shown in Fig. 2, where we display the Wilson coefficients as a function of the matching scale, varied in the interval $\mu_{M}\in[M_{\mathrm{OS}}/2,2M_{\mathrm{OS}}]$ and normalized to the value at $\mu_{M}=M_{\mathrm{OS}}$. In the case of $C_{duue}^{(-)}$, the small matching-scale dependence at LL results from an accidental numerical cancellation between $SU(3)_{c}$ and $U(1)_{Y}$ corrections, see Eq. (6.3).
@@
Figure 2: Comparison of the LL and NLL running of the Wilson coefficients due to the gauge interactions, normalized to the tree-level expression. The uncertainties are obtained by varying the matching scale between $\mu_{M}=M_{\mathrm{OS}}/2$ and $\mu_{M}=2M_{\mathrm{OS}}$. The solid lines include a resummation of large logarithms, while the dashed lines are the re-expanded LL and NLL results, truncated at NLO and NNLO, respectively.
In Fig. 2, we show the running of the Wilson coefficients due to the gauge interactions. The solid lines depict the numerical solutions of the RGEs, i.e., a full resummation of leading and next-to-leading logarithms. The dashed lines show a perturbative expansion of LL and NLL results to NLO and NNLO, respectively, as in Eqs. (6.3) and (6.3). As expected, for a large scale separation the resummation of large logarithms is indispensable. The NLL corrections are mostly of the order of the naive LL uncertainty estimate obtained from the matching-scale variation, with the notable exception of $C_{duue}^{(-)}$, where the scale variation significantly underestimates the NLL corrections due to the aforementioned numerical cancellation of $SU(3)_{c}$ and $U(1)_{Y}$ contributions. The NLL uncertainties estimated with the variation of the matching scale are so small that the error bands are not visible at the scale of the plots. Overall, the NLL corrections are perhaps surprisingly small, but we stress that in Fig. 2 we only show the evolution down to the electroweak scale. If we just consider QCD corrections, the single-gauge-coupling solution of the diagonalized NLL RGE has the simple form

### v2

@@ 1 Introduction
In these cases, BNV usually occurs at very high energies so that its effects are strongly suppressed at low energies, ensuring a sufficiently long proton lifetime to satisfy the experimental constraints. In such a setup, the BNV effects can be parametrized within effective field theories (EFTs) via higher-dimension operators, starting at dimension 6. As the BSM scale is usually much higher than the electroweak scale, one can use the SM Effective Field Theory (SMEFT) with explicit $SU(2)_{L}$ invariance to reduce the number of operators Buchmuller:1985jz ; Grzadkowski:2010es , see Refs. Crivellin:2023ter ; Beneito:2023xbk ; Gargalionis:2024nij ; Beneke:2024hox ; Gisbert:2024sjw ; Broussard:2025opd for recent work. For the evolution of the Wilson coefficients below the electroweak scale, the SMEFT should be matched to the low-energy EFT (LEFT). This EFT approach not only simplifies the calculations but also improves the precision since the potentially huge scale separation between the BNV scale and the hadronic scale mandates the resummation of large logarithms.
@@
The one-loop renormalization-group equations (RGEs) of these two EFTs at dimension 6 Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga ; Alonso:2014zka ; Jenkins:2017dyc , together with the tree-level matching of the SMEFT on LEFT Jenkins:2017jig can be used for computations at leading-logarithmic (LL) accuracy. For next-to-leading-log (NLL) accuracy the one-loop matching between SMEFT and LEFT Dekens:2019ept , the two-loop RGEs Buras:1989xd ; Buras:1991jm ; Buras:1992tc ; Ciuchini:1993vr ; Ciuchini:1993fk ; Buchalla:1995vs ; Chetyrkin:1997gb ; Buras:2000if ; Gorbahn:2016uoy ; Panico:2018hal ; deVries:2019nsu ; Bern:2020ikv ; Aebischer:2022anv ; Fuentes-Martin:2022vvu ; Aebischer:2023djt ; Jenkins:2023rtg ; Jenkins:2023bls ; Naterop:2023dek ; DiNoi:2023ygk ; Fuentes-Martin:2023ljp ; Morell:2024aml ; Aebischer:2024xnf ; Manohar:2024xbh ; DiNoi:2024ajj ; Born:2024mgz ; Naterop:2024cfx ; Fuentes-Martin:2024agf ; Aebischer:2025hsx ; Duhr:2025zqw ; Haisch:2025lvd ; Zhang:2025ywe ; Naterop:2025lzc ; Naterop:2025cwg ; Duhr:2025yor ; Haisch:2025vqj , as well as the calculation of all observables at next-to-leading order (NLO) are necessary. Furthermore, the one-loop matching of UV models on the SMEFT at NLO is needed, which has been automated to a large extent Carmona:2021xtq ; Fuentes-Martin:2022jrf ; Fuentes-Martin:2023ljp ; Guedes:2023azv ; Aebischer:2023nnv ; Thomsen:2024abg ; Guedes:2024vuf .
@@
For the BNV sector at dimension 6, the one-loop RGEs were calculated in Refs. Abbott:1980zj ; Alonso:2014zka ; Jenkins:2017dyc . Two- and three-loop QCD corrections were obtained in Refs. Nihei:1994tx ; Gracey:2012gx . In the LEFT, the two-loop RGEs are already available in multiple schemes Aebischer:2025hsx ; Naterop:2025lzc . [^1] The disagreement between these two references is expected to disappear with an upcoming revision of Ref. Aebischer:2025hsx . In the following, we rely on the results of Ref. Naterop:2025lzc . In the present article, we complete the two-loop RGE calculation for the BNV SMEFT operators at dimension 6. In addition to $SU(3)_{c}$ and $U(1)_{Y}$ gauge contributions as in the LEFT, our calculation includes the $SU(2)_{L}$ gauge and Yukawa interactions, as well as mixed gauge–Yukawa two-loop contributions and the Higgs self-interactions. In addition, we calculate the one-loop correction to the matching on SMEFT of the $S_{1}$ leptoquark, which can generate the Wilson coefficients of all four gauge-invariant operators. Therefore, our work completes the NLL EFT framework for the BNV sector at dimension 6, enabling refined analyses of the constraints from proton decay on heavy BNV new physics if combined with the NLO matching corrections Gracey:2012gx to non-perturbative determinations of the hadronic matrix elements Yoo:2021gql .
@@ 3.1 Dimensional regularization
As in the LEFT Naterop:2025lzc ; Aebischer:2025hsx , the two-loop renormalization of the BNV sector of the SMEFT does not involve any ill-defined $\gamma_{5}$-odd traces. Therefore, we perform the calculation in naive dimensional regularization (NDR) in $D=4-2\varepsilon$ space-time dimensions with anti-commuting $\gamma_{5}$, avoiding the issue of spurious symmetry breaking of the ’t Hooft and Veltman scheme and the need for finite symmetry-restoring renormalizations tHooft:1972tcz ; Breitenlohner:1977hr .
@@ 6.3 Cancellation of scheme and scale dependences
In the following, we check the cancellation of scheme and scale dependences between the two-loop RGEs and the one-loop matching. We start by defining the input parameters in the $\overline{\text{MS}}$ scheme at the fixed scale of the LQ mass $\mu=M$. These are the $\overline{\text{MS}}$ LQ mass $M(M)$ itself and the $\overline{\text{MS}}$ LQ Yukawa couplings $Y^{LL}(M)$, $Y^{RR}(M)$, $Y^{Q,LL}(M)$, and $Y^{Q,RR}(M)$. We then relate the $\overline{\text{MS}}$ parameters at $\mu=M$ to the ones at the matching scale $\mu=\mu_{M}$, to NLO or LL accuracy. Since $\mu_{M}\approx M$, this relation does not involve large logarithms and the difference between a fixed-order NLO relation and a resummation is immaterial. At one loop, the scale dependence of the $\overline{\text{MS}}$ LQ mass is
@@
whereas for the running of the LQ Yukawa couplings we find
@@
$\quad+Y^{Q,RR}\left(\langle Y^{RR}Y^{RR\dagger}\rangle+2\langle Y^{LL}Y^{LL\dagger}\rangle+2\langle Y^{Q,RR}Y^{Q,RR\dagger}\rangle+8\langle Y^{Q,LL}Y^{Q,LL\dagger}\rangle\right)\,.$ (6.7)
The corrections due to higher powers of the LQ Yukawa couplings are included for completeness but are not needed for the check of scheme- and scale-dependence cancellation.
Next, we determine the SMEFT Wilson coefficients at the scale $\mu=\mu_{M}$ using the one-loop matching. We subsequently apply the two-loop SMEFT RGEs to determine the Wilson coefficients at some low scale $\mu_{L}$, resumming large logarithms to NLL accuracy. The dependence on the matching scale is determined from the variation in the interval $\mu_{M}\in[M/2,2M]$.
For a phenomenological application, the renormalized $\overline{\text{MS}}$ input parameters should be related to physical observables. In the case of the LQ mass, the relation between $\overline{\text{MS}}$ mass $M$ at the scale $\mu$ and the pole mass $M_{\mathrm{OS}}$ is given by,
[…]
where $C_{F}=4/3$, $\alpha_{1}=g_{1}^{2}/(4\pi)$, $\alpha_{s}=g_{3}^{2}/(4\pi)$, $\langle\cdot\rangle$ denotes the trace in flavor space, and we neglect the fermion masses, i.e., we compute the pole mass in the unbroken phase. Note that due to the fermionic decay channels, the LQ acquires a width, i.e., the pole mass gets an imaginary part. We again included the effects due to higher powers of the LQ Yukawa couplings just for completeness. Similarly to the mass, the LQ Yukawa couplings should be related to the physical LQ decay widths at NLO, i.e., including virtual and real corrections.
For the illustration of the cancellation of scheme and matching-scale dependences, these relations to observables are not needed and we simply work with $\overline{\text{MS}}$ input parameters at the fixed scale $\mu=M$. For simplicity, we restrict the discussion to the gauge interactions, but with the full two-loop RGE results provided as supplementary material, the analysis can be extended to the Yukawa interactions. Using the flavor decompositions (2), which have diagonal gauge corrections, we can conveniently express the results in terms of the tree-level expressions Eqs. (6.1) and (6.1) with renormalized Yukawa couplings $Y(M)$ and $\overline{\text{MS}}$ mass $M(M)$.
For an analytic check of the matching-scale cancellation, we expand the solution of the RGEs in powers of the $\overline{\text{MS}}$ gauge couplings at $\mu=M$. At LL accuracy, re-expanded to NLO, this leads to
@@
$\frac{C_{duq\ell}(\mu_{L})}{C_{duq\ell}^{\mathrm{tree}}}\Big|^{\mathrm{NLL}}_{\mathrm{NNLO}}$ $=1-\left(\frac{11\alpha_{1}}{24\pi}+\frac{9\alpha_{2}}{8\pi}+\frac{\alpha_{s}}{\pi}\right)\log\!\left(\frac{\mu_{L}}{M}\right)$
$\quad+\left(-\frac{781\alpha_{1}^{2}}{1152\pi^{2}}+\frac{195\alpha_{2}^{2}}{128\pi^{2}}+\frac{9\alpha_{s}^{2}}{4\pi^{2}}+\frac{33\alpha_{1}\alpha_{2}}{64\pi^{2}}+\frac{11\alpha_{1}\alpha_{s}}{24\pi^{2}}+\frac{9\alpha_{2}\alpha_{s}}{8\pi^{2}}\right)\log^{2}\!\left(\frac{\mu_{L}}{M}\right)$
$\quad-\frac{(1+2a_{\mathrm{ev}})\alpha_{1}}{12\pi}$
$\quad+\bigg(\begin{aligned} &\frac{(811-1136a_{\mathrm{ev}})\alpha_{1}^{2}}{2304\pi^{2}}-\frac{541\alpha_{2}^{2}}{256\pi^{2}}-\frac{31\alpha_{s}^{2}}{24\pi^{2}}\\ &+\frac{3(5+8a_{\mathrm{ev}})\alpha_{1}\alpha_{2}}{128\pi^{2}}+\frac{(37+24a_{\mathrm{ev}})\alpha_{1}\alpha_{s}}{144\pi^{2}}+\frac{9\alpha_{2}\alpha_{s}}{16\pi^{2}}\bigg)\log\!\left(\frac{\mu_{L}}{M}\right)\end{aligned}$
$\quad\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}+\left(-\frac{1379\alpha_{1}^{2}}{2304\pi^{2}}+\frac{541\alpha_{2}^{2}}{256\pi^{2}}+\frac{31\alpha_{s}^{2}}{24\pi^{2}}-\frac{3\alpha_{1}\alpha_{2}}{128\pi^{2}}-\frac{25\alpha_{1}\alpha_{s}}{144\pi^{2}}-\frac{9\alpha_{2}\alpha_{s}}{16\pi^{2}}\right)\log\!\left(\frac{\mu_{M}}{M}\right)\,$, (6.10)
where the couplings are evaluated at $\mu=M$. The NLL result allows us to make a few observations. The first two lines in Eq. (6.3) contain the LL terms, while the next three lines are NLL terms (both truncated at NNLO). While the leading logs are scheme independent, the NLL terms contain scheme dependences, in particular a dependence on $a_{\mathrm{ev}}$. The reason is that the Wilson coefficient itself is a scheme-dependent quantity beyond tree-level. The last line contains a logarithm of the matching scale over the LQ mass, which is not a large logarithm. These terms (in gray) are beyond the accuracy of our calculation, as we did not compute the renormalization of the LQ model at two loops. We find that the dependence on the matching scale $\mu_{M}$ is shifted to NNLO as expected, i.e., all NLO matching-scale dependences cancel. Regarding the scheme dependence, we find that the matching-scale-dependent terms do not depend on $a_{\mathrm{ev}}$: we observe a cancellation of such $a_{\mathrm{ev}}$-dependent terms between the two-loop RGE and the one-loop matching. The remaining $a_{\mathrm{ev}}$-dependence in Eq. (6.3) will cancel instead at the low scale $\mu_{L}$, when combining the two-loop RGE with a NLO matrix-element or matching calculation. The same structure is found for the NLL results of the other Wilson coefficients, which are included in the supplementary material.
@@
$C(\mu=M,\mu_{M})/C(\mu=M,\mu_{M}=M)$
$C_{duq\ell}LLNLL$$C_{qque}$$C_{duue}^{(+)}LLNLL$$C_{duue}^{(-)}$$\mu_{M}/M$$S_{qqq\ell}LLNLL$$M_{qqq\ell}$
Figure 1: Comparison of the LL and NLL dependence of the Wilson coefficients on the matching scale $\mu_{M}$, normalized to a matching scale of $\mu_{M}=M$, which is chosen at the unification scale, $M=6.5\times 10^{15}\,\text{GeV}$. We only include the contribution of the gauge interactions.
The reduction of the matching-scale dependence from NLO to NNLO is shown in Fig. 1, where we display the Wilson coefficients as a function of the matching scale, varied in the interval $\mu_{M}\in[M/2,2M]$ and normalized to the value at $\mu_{M}=M$. In the case of $C_{duue}^{(-)}$, the small matching-scale dependence at LL results from an accidental numerical cancellation between $SU(3)_{c}$ and $U(1)_{Y}$ corrections, see Eq. (6.3).
@@
Figure 2: Comparison of the LL and NLL running of the Wilson coefficients due to the gauge interactions, normalized to the tree-level expression. The uncertainties are obtained by varying the matching scale between $\mu_{M}=M/2$ and $\mu_{M}=2M$. The solid lines include a resummation of large logarithms, while the dashed lines are the re-expanded LL and NLL results, truncated at NLO and NNLO, respectively.
In Fig. 2, we show the running of the Wilson coefficients due to the gauge interactions. The solid lines depict the numerical solutions of the RGEs, i.e., a full resummation of leading and next-to-leading logarithms. The dashed lines show a perturbative expansion of LL and NLL results to NLO and NNLO, respectively, as in Eqs. (6.3) and (6.3). As expected, for a large scale separation the resummation of large logarithms is indispensable. The NLL corrections are mostly of the order of the naive LL uncertainty estimate obtained from the matching-scale variation, with the exception of $C_{duue}^{(-)}$, where the scale variation significantly underestimates the NLL corrections due to the aforementioned numerical cancellation of $SU(3)_{c}$ and $U(1)_{Y}$ contributions. The NLL uncertainties estimated with the variation of the matching scale are so small that the error bands are not visible at the scale of the plots. Additional finite one-loop corrections arise when relating the $\overline{\text{MS}}$ parameters to observables, which are not included in Fig. 2, see Eq. (6.8). Overall, the NLL corrections are surprisingly small, but we stress that in Fig. 2 we only show the evolution down to the electroweak scale. If we just consider QCD corrections, the single-gauge-coupling solution of the diagonalized NLL RGE has the simple form